determine whether the given ordered pair is a solution
(7/25, -1/25)
4x+3y=1

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

determine whether the given ordered pair is a solution
(7/25, -1/25)
4x+3y=1
------------------
Substitute to see if you get a true statement:
4(7/25)+3(-1/25) = 1
-----
(28 - 3)/25 = 1
-----
25/25 = 1
---
1 = 1
----
The given ordered pair is a solution.
===============================================
Cheers,
Stan H.
===============

Answer by richwmiller(17219)

4*7=28
28-3=25
25/25=1
yes